Margo has 40 millimeters of 25% solution. how many millimeters of 60% solution should she add to obtain the required 30% solution?
Answer by stanbon(75887)

Margo has 40 millimeters of 25% solution. how many millimeters of 60% solution should she add to obtain the required 30% solution?
----
Equation:
active + active = active
0.25*40 + 0.60x = 0.30(40+x)
----
Multiply thru by 100 to get:
25*40 + 60x = 30*40 + 30x
30x = 5*40
----
x = (4/3)5 = 20/3 = 6 2/3 ml (amt of 60% solution needed)
==================================
Cheers,
Stan H.
===============